# Graph-based RAG Enhancement via Global Query Disambiguation and Dependency-Aware Reranking

### v1

@@ Abstract
Contemporary graph-based retrieval-augmented generation (RAG) methods typically begin by extracting entities from user queries and then leverage pre-constructed knowledge graphs to retrieve related relationships and metadata. However, this pipeline’s exclusive reliance on entity-level extraction can lead to the misinterpretation or omission of latent yet critical information and relations. As a result, retrieved content may be irrelevant or contradictory, and essential knowledge may be excluded, exacerbating hallucination risks and degrading the fidelity of generated responses. To address these limitations, we introduce PankRAG, a framework that combines a globally aware, hierarchical query-resolution strategy with a novel dependency-aware reranking mechanism. PankRAG first constructs a multi-level resolution path that captures both parallel and sequential interdependencies within a query, guiding large language models (LLMs) through structured reasoning. It then applies its dependency-aware reranker to exploit the dependency structure among resolved sub-questions, enriching and validating retrieval results for subsequent sub-questions. Empirical evaluations demonstrate that PankRAG consistently outperforms state-of-the-art approaches across multiple benchmarks, underscoring its robustness and generalizability.
Graph-based RAG Enhancement via Global Query Disambiguation and Dependency-Aware Reranking
Ningyuan Li1, Junrui Liu1, Yi Shan1 , Minghui Huang1  and Tong Li1 1 College of Computer Science Beijing university of technology ningy@emails.bjut.edu.cn, liujunrui@emails.bjut.edu.cn,litong@bjut.edu.cn
@@ 1 Introduction
Large language models (LLMs) exhibit exceptional performance across diverse tasks by leveraging pre-training on massive text corpora to generate appropriate responses. While LLMs’ capabilities can be further enhanced through fine-tuning on downstream tasks, finite parameter capacity limits their ability to store comprehensive and evolving world knowledge. Retrieval-augmented generation (RAG)  Lewis et al. (2020); Guu et al. (2020); Gao et al. (2023); Ram et al. (2023); Fan et al. (2024); Izacard et al. (2023); Gao et al. (2024)has emerged as the predominant solution, owing to its straightforward implementation. It enables static models to access up-to-date external knowledge while preserving interpretability and provenance tracking.
@@
Retrieval-augmented generation (RAG) methods dynamically integrate heterogeneous knowledge bases Sun et al. (2023); Edge et al. (2024); Guo et al. (2024); Ma et al. (2024) to enhance contextual reasoning and factual grounding. Among these, knowledge graphs  Chen et al. (2020); Jin et al. (2024) serve as structured frameworks that encode information through interconnected entities and relationships, explicitly modeling semantic dependencies to enable human-like understanding and multi-hop logical inference. Recently, many approaches have integrated RAG with knowledge graphs. For example, GraphRAG Edge et al. (2024) enhances private-text question answering by precomputing community summaries for clusters of related entities, dynamically synthesizing these summaries into contextually unified responses during query processing. LightRAG  Guo et al. (2024) employs a graph-text hybrid architecture that embeds structural relationships into document indexing and retrieval. Its hierarchical retrieval system combines graph-based relational reasoning with vector similarity matching, enabling efficient knowledge extraction while preserving semantic fidelity in multi-hop queries.
However, current graph-based RAG methods exhibit critical limitations of inaccuracies in the interpretation of critical information or relationships when handling specific and abstract complex queries. Specific complex queries (SCQ) demand entity-centric granularity, requiring precise retrieval of discrete entities and their relational contexts within structured knowledge systems, e.g., "In which direction would you travel to reach Southampton from the city connected to it by rail?". For specific queries, existing methods fail to capture latent information embedded in multi-hop relationships. For example, when GraphRAG’s local search retrieves content for the prior specific query, it returns extensive information about “Southampton,” most of which, beyond the city’s geographic location and its rail-connected neighbor, is irrelevant and may exacerbate LLM hallucinations. Abstract complex queries(ACQ) emphasize conceptual synthesis beyond entity-level precision, addressing domain-wide implications or systemic interactions, e.g., “How can theoretical principles from programming language theory be made accessible to new developers?”,(A detailed case study analyzing this example ACQ is presented in subsection A.3). For ACQ, methods often overlook the implicit context underlying conceptual abstractions. When LightRAG retrieves content for the prior ACQ, it prioritizes formal syntax definitions, while teaching tools, like step-by-step tutorials, error-driven case studies, and diverse code examples—provides richer contextual guidance, enabling the LLM to generate more in-depth responses. These limitations underscore the insufficiency of single-phase, query-as-is retrieval paradigms in addressing real-world SCQ and ACQ, challenging current graph-based RAG methods.
@@
In this paper, we propose PankRAG, which resolves queries through a hierarchical resolution path with global awareness, augmented by a dependency-aware reranking mechanism. Specifically, our approach leverages the robust planning capabilities of LLMs to generate globally-aware (i.e., holistically planned rather than stepwise) problem-solving pathways, explicitly revealing parallel and progressive relationships among sub-questions. Next, we refine complex queries by integrating the implicit critical information resolved from prior sub-question resolutions. This dual-level process produces a hierarchical resolution pathway with global awareness, enhancing interpretability through contextual grounding. It enables LLMs to dynamically disambiguate and contextualize ambiguous or polysemous queries before final answer generation. During execution, our dependency-aware reranking mechanism augments and reorders retrieved content for sub-questions dependent on prior resolutions by integrating their resolved dependencies. This process refines the retrieved context by integrating resolved sub-question dependencies, thereby enhancing the alignment between retrieved contexts and dependency-driven sub-question answers challenging for LLMs to capture and resolve.
We evaluate PankRAG on four benchmark datasets for ACQ and two for SCQ, demonstrating that it consistently outperforms state-of-the-art baselines across six benchmarks. Compared to NaiveRAG, GraphRAG, and LightRAG, PankRAG achieves average win rates of 87.0%, 59.78% and 61.59% respectively in ACQ datasets, and achieves average metric improvements of 17.42%, 11.61%, 11.08% in SCQ datasets.
@@
  • We propose a hierarchical query decomposition strategy that systematically plans globally-aware problem-solving pathways, enhancing LLMs’ capacity to capture and resolve implicit information and relationships, fostering alignment between query intent and generated responses.
@@
  • We propose a dependency-aware reranking mechanism that augments and reorders retrieved content for the current sub-question by integrating resolved sub-question dependencies.
@@
  • PankRAG demonstrates consistent superiority over state-of-the-art baselines, achieving comprehensive improvements on four ACQ benchmarks evaluated across six dimensions and two SCQ benchmarks assessed via four metrics.
Figure 1: The overall workflow of PankRAG. During Globally-Aware Planning, queries are decomposed into colored resolution pathways. In the bottom-up phase, two parallel sub-question sequences are generated: a blue sequence for interdependent sub-questions that require sequential processing, and an green independent sequence that can be resolved concurrently. In the top-down phase, an orange sequence rewrites complex queries, incorporating resolved sub-question dependencies to enhance clarity.
2 Releated Works
2.1 Retrieval-Augmented Generation Based On Chunks
RAG Fan et al. (2024); Gao et al. (2024) is widely used to address the timeliness of knowledge bases and hallucination issues Yao et al. (2023) in large language models (LLMs). Additionally, this approach is more cost-effective and easier to implement compared to methods like fine-tuning or retraining, significantly reducing computational and time resources. NaiveRAG  Gao et al. (2024) is a method that typically processes documents and queries by splitting them into smaller text chunks, converting these chunks into vector representations. The retrieved chunks, which are selected by comparing similarity with queries in vector space, are then used to generate responses through LLMs.
Numerous advanced RAG methods, such as MemoRAG Qian et al. (2024), RQ-RAG Chan et al. (2024), and others Asai et al. (2023); Lin et al. (2023), build upon the foundational framework of NaiveRAG, introducing specialized improvements. MemoRAG  Qian et al. (2024) is an advanced RAG framework empowered by global memory-augmented retrieval. It combines two specialized components: a lightweight system that generates draft answers as "global memory" to guide precise retrieval, and a powerful system that crafts the final answer using the retrieved information. RQ-RAG is a method designed to tackle ambiguous or complex queries. The model is equipped with explicit capabilities for query rewriting, decomposition, and disambiguation to enhance retrieval accuracy. Inspired by two methods, our proposed method integrates query decomposition and query rewriting principles, enhancing comprehension of hidden critical information or relationships within queries.
However, most text-based RAG methods utilize a flat structured database, failing to capture critical inter-chunk relationships, which constrains the quality of LLM-generated responses. While advanced RAG variants improve generation accuracy, their reliance on retrieving a limited subset of documents restricts comprehensive global information integration. It ultimately hinders their ability to address tasks requiring full-context understanding, such as ACQ.
2.2 Retrieval-Augmented Generation Based On KGs
Apart from a database of texts, a knowledge graph Jin et al. (2024); Chen et al. (2020), a structured framework organizing knowledge through interconnected entities and relationships, is increasingly employed as a valuable source for RAG methods. Unlike traditional RAG Gao et al. (2024), Graph-based RAG methods Sun et al. (2023); Ma et al. (2024)leverage graphs either as knowledge carriers or indexing tools to enable efficient retrieval. By modeling relationships between nodes, they facilitate the discovery of topic-centered or entity-centered knowledge pieces, ensuring thorough and contextually coherent retrieval. Furthermore, the graph structure contained dependencies or relationships between entities, which enhances both precision and scalability compared to flat databases.
Graph-based RAG methods differ in how they leverage the underlying graph data. GraphRAG  Edge et al. (2024) is designed to enhance question answering over private text corpora by using a knowledge graph. It pre-generates community summaries for clusters of related entities and dynamically aggregates them into coherent final responses during query resolution. LightRAG  Guo et al. (2024) is a novel framework that integrates graph structures into text indexing and retrieval. Employing a dual-level retrieval mechanism, it enhances comprehensive knowledge discovery and accelerates response generation without compromising contextual accuracy.
While prior research introduces knowledge graphs as a structured database, inaccuracies in the interpretation of hidden critical information or relationships within queries remain persistent challenges. Specifically, while knowledge graphs provide RAG systems with robust knowledge representation and enhanced retrieval performance, their practical effectiveness depends on accurately comprehending latent critical information and relationships within queries to identify precise retrieval targets. Our proposed framework is specifically designed to address this challenge. We also notice a concurrent work, CogGRAG Cheng et al. (2025) that leverages the principle of query decomposition. While they predominantly focus on problem decomposition strategy, our research prioritizes the acquisition of problem-solving pathways, with particular emphasis on the interplay between parallel interrelationships within these pathways and the impact of resolved sub-question dependencies on contextualizing information retrieved for the current sub-question.
@@ 3 Methodology
In this section, we detail our proposed framework, which comprises three core components. First, Graph Construction extracts entities and relationships from external knowledge sources to build a structured knowledge graph. Next, Globally-Aware Planning formulates a query resolution pathway by identifying latent critical information and relationships, improving LLMs’ comprehensiveness of complex queries. Finally, during Execution, the LLM processes sub-questions either concurrently or sequentially based on their interrelationships. After retrieving context for a sub-question, the dependency-aware reranking mechanism augments and validates the content when the current sub-question depends on a prior sub-question. The LLM then leverages this refined context to generate accurate responses. The overall workflow of FG-RAG is shown in Figure 1.
@@ 3.1 Graph Contruction
The source document corpus $D$ is segmented into $n$ text chunks $\{D_{i}\}^{n}_{i=1}$. As the Equation 1, for each chunk $D_{i}$, the LLM performs entity recognition and relationship extraction, abstracting structured semantic elements including entities $V_{i}$ and relations $E_{i}$ from unstructured text via prompting strategies. A knowledge graph $G$ is then constructed by instantiating nodes representing entities $\{V_{i}\}$ and edges representing relationships $\{E_{i}\}$ in Equation 2.
@@
Next, Multiple communities $\{Com_{i}\}$ are aggregated into a higher-level community $H_{i}$ via Leiden clustering, and this process is iteratively applied to hierarchically construct a multi-layered community structure. Finally, multi-scale community summarization is executed by the LLM, which populates a summary template with various elements (nodes, edges, and associated claims).
@@ 3.2 Globally-Aware Planning
Figure 2: Operation example of Globally-Aware Planning. The color assignments follow the configuration illustrated in the caption of Figure 1. [A1.1], [A1.2], and [A2.1] correspond to the resolved outputs of sub-questions SubQ[1.1], SubQ[1.2], and SubQ[2.1], respectively.
We propose a hierarchical query decomposition strategy that synergizes bottom-up relationship decomposition with top-down contextual rephrasing to address both implicit parallel or progress relationships and ambiguous or polysemous expressions in complex queries. Unlike conventional iterative approaches (e.g., Chain-of-Thought, CoT), which suffer from fixed iteration counts and myopic decomposition prone to local optima, our strategy dynamically adjusts the granularity and sequence of reasoning steps based on query complexity.
In the bottom-up phase, two crucial relationship types are detected. First, parallelizable task relationships are identified to enable concurrent resolution of mutually independent sub-questions, optimizing computational efficiency by reducing redundant resource contention and minimizing latency. Second, latent dependencies are systematically extracted to ensure prerequisite sub-questions are solved before dependent tasks.
Simultaneously, the top-down phase disambiguates polysemous expressions into logically coherent substructures, combining the resolved sub-question dependencies. This dual mechanism reduces the risk of deviation in reasoning paths and enhances computational efficiency by eliminating redundant iterations, achieving a balanced trade-off between precision and scalability. The detailed workflow is presented below, and a real-world query in Multihop-RAG dataset example is above in Figure 2.
  • Bottom-up phase: In the Parallel Relationship Detection step, parallel semantic relationships—such as independent sub-questions or co-occurring entities—are identified. If no parallelism is detected, a single resolution sequence is initialized. When parallelism exists, multiple concurrent sequences (SubQ1.x and SubQ2.x in Figure 2) are adaptively spawned to optimize concurrent execution. Concurrently, latent progress relationships are extracted from the query. Upon identifying distinct relationships within the query, sub-questions are generated and formalized into a directed acyclic graph (DAG), where nodes represent sub-questions and edges encode their relationships. Finally, topological sorting is applied to enforce prerequisite-resolution order, ensuring logical coherence during execution.
@@ 3.3 Execution
3.3.1 Retrieval
Following GraphRAG, we apply distinct retrieval strategies for different query types.
For SCQ, we use a local search approach that employs vector similarity metrics to identify nodes explicitly referenced in the query, then expands to neighboring nodes to retrieve related relationships and community context. This method ensures precise grounding in entity-centric evidence while maintaining low computational overhead.
For ACQ, we perform hierarchical community retrieval, dynamically selecting between leaf-level summaries for detailed information and higher-level summaries for thematic overviews. This multi-scale strategy balances breadth and depth to search for relevant information for complex queries.
Table 1: Performance across four datasets and six evaluation dimensions in terms of averaged win rates.
@@ 3.3.2 Dependency-Aware Reranking Mechanism
The dependency-aware reranking mechanism is triggered when resolving a sub-question that depends on previously resolved sub-questions. It dynamically adjusts retrieval priorities by indirectly leveraging these resolved dependencies. Unlike conventional methods that directly incorporate resolved sub-question dependencies, risking cumulative hallucination errors across iterative reasoning steps, our mechanism mitigates such propagation by reinforcing contextually relevant content and filtering out incongruent evidence through semantic similarity validation. If no dependencies are present, the generation proceeds without reranking, preserving computational efficiency while maintaining robustness against cascading inaccuracies.
The reranking mechanism computes a composite score by weighted summation of two components: the intrinsic quality score $R_{i}$ of retrieved content and the semantic similarity $M_{i}$ between retrieved content and resolved dependent sub-question answers. The intrinsic quality score $R_{i}$ of retrieved content is generated and retained during the original retrieval phase Edge et al. (2024). The semantic similarity $M_{i}$ is computed as the cosine similarity between resolved sub-question dependencies $\{sub_{i}\}$ and the initially retrieved text chunks $\{ori_{i}\}$. Each chunk’s similarity score is updated and then linearly combined with its original retrieval quality score to optimize prioritization. The weight parameters $\alpha$ (quality) and $\beta$ (relevance) were selected by testing on the dataset(subsection 4.4), ensuring maximal retrieval precision while mitigating hallucination propagation. This process can be formally described as follows.
@@
Combined $\text{Combined Score}=\alpha\times\mathcal{R}_{i}+\beta\times\mathcal{M}_{i}$ (5)

### v2

# PankRAG: Enhancing Graph Retrieval via Globally Aware Query Resolution and Dependency-Aware Reranking Mechanism

@@ Abstract
Recent graph-based RAG approaches leverage knowledge graphs by extracting entities from a query to fetch their associated relationships and metadata. However, relying solely on entity extraction often results in the misinterpretation or omission of latent critical information and relationships. This can lead to the retrieval of irrelevant or contradictory content, as well as the exclusion of essential information, thereby increasing hallucination risks and undermining the quality of generated responses. In this paper, we propose PankRAG, a framework designed to capture and resolve the latent relationships within complex queries that prior methods overlook. It achieves this through a synergistic combination of a globally-aware hierarchical resolution pathway and a dependency-aware reranking mechanism. PankRAG first generates a globally aware resolution pathway that captures parallel and progress relationships, guiding LLMs to resolve queries through a hierarchical reasoning path. Additionally, its dependency-aware reranking mechanism utilizes resolved sub-question dependencies to augment and validate the retrieved content of the current unresolved sub-question. Experimental results demonstrate that PankRAG consistently outperforms existing state-of-the-art methods, underscoring its generalizability. The code is available at:https://github.com/Ninggggy/PankRAG.
Index Terms—  Question answering, Retrieval-augmented generation, Query decomposition.
@@ 1 Introduction
Large language models (LLMs)  [17, 18]exhibit exceptional performance across diverse tasks by leveraging pre-training [11] on massive text corpora to generate appropriate responses. While LLMs’ capabilities can be further enhanced through fine-tuning [2] on downstream tasks, finite parameter capacity limits their ability to store comprehensive and evolving world knowledge. Retrieval-augmented generation (RAG)  [15, 9, 6, 20, 5, 12, 7, 14]has emerged as the predominant solution, owing to its straightforward implementation. It enables static models to access up-to-date external knowledge while preserving interpretability and provenance tracking.
@@
RAG dynamically integrate heterogeneous knowledge bases [21, 3, 8, 16, 13] to enhance contextual reasoning and factual grounding. Among these, knowledge graphs  [1] serve as structured frameworks that encode information through interconnected entities and relationships, explicitly modeling semantic dependencies to enable human-like understanding and multi-hop logical inference. Recently, many approaches have integrated RAG with knowledge graphs to enhance factual grounding and multi-hop inference, with prominent examples including GraphRAG[3], LightRAG[8], HippoRAG[13], and G-Retriever[10]. For instance, GraphRAG[3] enhances question answering by adaptively synthesizing precomputed summaries of entity communities. LightRAG[8] leverages a hierarchical retrieval system that fuses graph-based reasoning with vector matching to ensure efficient knowledge extraction and semantic fidelity for multi-hop queries. However, the single-phase retrieval approach of current graph-based RAG methods is ill-suited for deep reasoning, causing them to misinterpret or omit latent critical information and relationships when handling complex queries. For Specific Complex Queries (SCQ) [8] requiring precise multi-hop inference, they often retrieve broad, irrelevant information centered on a core entity instead of the specific relational facts, increasing the risk of LLMs’ hallucinations. Similarly, for Abstract Complex Queries (ACQ) [8] demanding conceptual synthesis, they tend to fetch superficial definitions while overlooking the rich, implicit context required for a nuanced response.
@@
In this paper, we propose PankRAG, which resolves complex queries through a hierarchical reasoning pathway with global awareness, augmented by a dependency-aware reranking mechanism. Specifically, our approach leverages the robust planning capabilities of LLMs to generate a globally-aware (i.e., holistically planned rather than stepwise) resolution pathway, explicitly revealing parallel and progress relationships among sub-questions. Next, we refine complex queries by integrating the implicit critical information resolved from prior sub-question resolutions. This dual-level process produces a hierarchical resolution pathway with global awareness, significantly enhancing retrieval performance and response quality. It enables LLMs to dynamically disambiguate and contextualize ambiguous or polysemous queries before final answer generation. During execution, our dependency-aware reranking mechanism augments and validates retrieved content for sub-questions. This process refines the retrieved context by integrating resolved sub-question dependencies, thereby enhancing the alignment between retrieved contexts and resolved sub-question dependencies.
@@
  • We propose a hierarchical query decomposition strategy that generates a globally-aware resolution pathway, enabling LLMs to capture and resolve latent information and relationships, ensuring a tight alignment between generated responses and query intent.
@@
  • We propose a dependency-aware reranking mechanism that augments and validates retrieved content for the current sub-question by integrating resolved sub-question dependencies.
@@
  • PankRAG demonstrates consistent superiority over state-of-the-art baselines, achieving comprehensive improvements on four ACQ datasets evaluated across six dimensions and two SCQ datasets assessed via four metrics.
Fig. 1: The overall workflow of PankRAG.
@@ 2.1 Graph Contruction
The source document corpus $D$ is segmented into $n$ text chunks $\{D_{i}\}^{n}_{i=1}$. As the Equation 1, for each chunk $D_{i}$, the LLM performs entity recognition and relationship extraction, abstracting entities $V_{i}$ and relations $E_{i}$ from unstructured text via prompting strategies. A knowledge graph $G$ is then constructed by instantiating nodes representing entities $\{V_{i}\}$ and edges representing relationships $\{E_{i}\}$ in Equation 2.
@@
Next, Multiple communities $\{Com_{i}\}$ are aggregated into a higher-level community $H_{i}$ via Leiden clustering, and this process is iteratively applied to hierarchically construct a multi-layered community structure. Finally, multi‑scale community summarization is executed by the LLM.
@@ 2.2 Globally-Aware Planning
Fig. 2: Operation example of Globally-Aware Planning. [A1.1], [A1.2], and [A2.1] correspond to the resolved outputs of SubQ[1.1], SubQ[1.2], and SubQ[2.1], respectively.
We propose a hierarchical query decomposition strategy that synergizes bottom-up relationship decomposition with top-down contextual rephrasing to address both implicit parallel or progress relationships and ambiguous or polysemous expressions in complex queries. Unlike conventional iterative approaches (e.g., Chain-of-Thought, CoT), which suffer from fixed iteration counts and myopic decomposition prone to local optima, our strategy dynamically adjusts the granularity and sequence of reasoning steps based on the query’s inherent structure.
  • Bottom-up phase: In the Parallel Relationship Detection step, parallel semantic relationships—such as independent sub‑questions or co‑occurring entities—are identified to enable parallel execution, thereby optimizing computational efficiency by minimizing resource contention. If no parallelism is detected, a single resolution sequence is initialized. When parallelism exists, multiple concurrent sequences (SubQ1.x and SubQ2.x in Figure 2) are adaptively spawned to optimize concurrent execution. Concurrently, latent progress relationships are extracted from the query. Upon identifying distinct relationships within the query, sub-questions are generated and formalized into a directed acyclic graph (DAG), where nodes represent sub‑questions and edges encode their relationships. Finally, topological sorting is applied to enforce prerequisite‑resolution order, ensuring logical coherence during execution.
@@
This dual mechanism reduces the risk of deviation in reasoning paths and enhances computational efficiency by eliminating redundant iterations, achieving a balanced trade‑off between precision and scalability. The detailed workflow is above inFigure 1, and a real-world query in the Multihop-RAG [22] dataset example is presented in Figure 2.
@@ 2.3 Execution
Following GraphRAG[3], we tailor our retrieval strategy to the query type. For SCQs, we employ a local search that targets query-referenced nodes and expands to their neighbors, ensuring precise entity-centric grounding. For ACQs, we utilize hierarchical community retrieval, dynamically selecting between detailed leaf-level and thematic high-level summaries to balance retrieval breadth and depth.
Table 1: Overall Performance on the ACQ Datasets.
@@
HippoRAG PankRAG HippoRAG PankRAG HippoRAG PankRAG HippoRAG PankRAG
Comprehensiveness 45.6% 54.4% 48.0% 52.0% 42.8% 57.2% 46.8% 53.2%
Diversity 42.8% 57.2% 43.2% 56.8% 36.0% 74.0% 44.0% 56.0%
logicality 31.2% 68.8% 43.6% 56.4% 42.8% 57.2% 44.4% 55.6%
relevance 40.4% 59.6% 41.2% 58.8% 47.2% 52.8% 44.8% 55.2%
coherence 32.4% 67.6% 42.4% 57.6% 39.6% 60.4% 45.2% 54.8%
Empowerment 46.0% 54.0% 46.8% 53.2% 40.8% 59.2% 41.2% 58.8%
@@ 2.3.1 Dependency-Aware Reranking Mechanism
The dependency‑aware reranking mechanism is triggered when resolving a sub‑question that depends on previously resolved sub‑questions. It dynamically adjusts retrieval priorities by indirectly leveraging these resolved dependencies. Unlike conventional methods that directly incorporate resolved sub-question dependencies, risking cumulative hallucination errors across iterative reasoning steps, our mechanism mitigates such propagation by reinforcing contextually relevant content and filtering out incongruent evidence through semantic similarity validation. If no dependencies are present, the generation proceeds without reranking, preserving computational efficiency while maintaining robustness against cascading inaccuracies.
The reranking mechanism computes a composite score by weighted summation of two components: the intrinsic quality score $R_{i}$ of retrieved content and the semantic similarity $M_{i}$ between retrieved content and resolved dependent sub-question answers.
@@
The intrinsic quality score $R_{i}$ of retrieved content is generated and retained during the original retrieval phase [3]. The semantic similarity $M_{i}$ is computed as the cosine similarity between resolved sub-question dependencies $\{sub_{i}\}$ and the initially retrieved text chunks $\{ori_{i}\}$. Each chunk’s similarity score is updated and then linearly combined with its original retrieval quality score to optimize prioritization. The weight parameters $\alpha$ (quality) and $\beta$ (relevance) were selected by testing on the dataset(subsection 3.4), ensuring maximal retrieval precision while mitigating hallucination propagation. This process can be formally described as follows.
$\text{Combined Score}=\alpha\times\mathcal{R}_{i}+\beta\times\mathcal{M}_{i}$ (5)
2.3.2 Generation
$Ans=\textit{LLM}(Q,(D,K))$ (6)
The generation process integrates resolved dependencies $D$ and dependency-enhanced reranking knowledge $K$ into a consolidated context, which is utilized by the LLM to generate reliable responses.